Consider the following all-integer linear program: Max x1+x2

s.t 4x1+6x2 <= 22

x1+5x2<= 15

2x1+x2<=9

x1,x2>=0 integer

Solve in Excel Solver and AMPL.
Answer #1

Objective Function

Max Z = x1+x2

Constraints

4x1+6x2 <= 22

x1+5x2 <= 15

2x1+x2 <=9

x1,x2 >= 0 integer

Configure in Excel as below -

The blue cells needs to be optimized to get the max value in green cell

Run solver to get